The question ask is to "solve by factoring" the problem listed below. The answer given is -1, -12.
x^2-13x=-12
x^2-13x+12 = -12+12
x^2-13x+12 = 0
I'm stuck past this point!

Answer by stanbon(75887)

The question ask is to "solve by factoring" the problem listed below. The answer given is -1, -12.
x^2-13x=-12
x^2-13x+12 = -12+12
x^2-13x+12 = 0
(x-12)(x-1)= 0
x = 12 or x = 1
============
Cheers,
Stan H.